OpenStudy (anonymous):

A cylinder of diameter 10cm and a height of 20 cm is filled with water . if a sphere of diameter 10 cm is completely submerged inside the cylinder ,what will be the increase in the level of water, in the cylinder?

OpenStudy (anonymous):

1) Find volume of the sphere 2) Equate the volume of sphere with formula of volume of cylinder to find increase in height

OpenStudy (anonymous):

V of sphere= (4/3) r^3 =(4/3) 5^3 =500/3

OpenStudy (anonymous):

V of cyl=pi r^2h =pi(5^2) (20) =500pi

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (shubhamsrg):

Volume of cylinder = pir^2 h = 500pi volume of sphere immersed = 4/3pi r^3 = 500pi/3 New volume of cylinder = Initial cylinder volume + Sphere volume. pi*25*h = 500pi +500pi/3 h = 20(4/3) =80/3 hence change in height = h-20 = 20/3

OpenStudy (anonymous):

Thank u @shubhamsrg